# GAAPO: Genetic Algorithmic Applied to Prompt Optimization

### v1

@@ 3 Data
  • The MMLU-Pro (Massive Multitask Language Understanding Professional)[24] dataset extends the famous MMLU [25] dataset by complexifying it to a professional level, with our focus on two key subcategories. The Engineering subcategory evaluates technical understanding across various engineering disciplines, testing knowledge of fundamental principles, technical specifications, and complex problem-solving approaches encountered in professional practice. The Business subcategory assesses comprehension of management principles, corporate strategy, financial decision-making, and organizational behavior through practical business scenarios.
@@ 4.1.1 GAAPO: Genetic Algorithmic Applied to Prompt Optimization
The genetic algorithm framework incorporates multiple prompt generation methods, each implementing distinct optimization strategies as detailed in sections 4.1.2 and 4.1.3. These methods, summarized in Table 1, represent diverse approaches to prompt optimization, each with its own strengths and limitations. The hybrid nature of GAAPO leverages this diversity by combining these complementary strategies within a single optimization framework. This integration enables the algorithm to capitalize on the advantages of each method while mitigating their individual limitations through iterative application of varied optimization approaches. The synergistic combination of these methods allows for more comprehensive exploration of the prompt space than would be possible with any single strategy.
@@ 4.1.3 Generation methods: random evolutions
Each mutation creates a new variant of the original prompt (examples are presented in annex 6.0.1, potentially discovering more effective formulations. Like genetic mutations in nature, these modifications can range from subtle adjustments to significant transformations, allowing for both local and global exploration of the prompt space. This random but structured approach enables the discovery of novel prompt variations that might not be obvious through deterministic methods.
@@ 4.3.3 Evaluation Process
  • Initial evaluation on validation set to establish baseline performance
  • Generational evaluation to select promising candidates
  • Final testing on the held-out test set to measure true generalization
@@ 4.3.4 Metrics
Metric used for the multi classification task of the ETHOS dataset is a strict accuracy: a prediction is considered correct only when all labels for a given sample are correctly identified.
@@ 4.4 Experiments
300 samples were extracted to the orignal ETHOS dataset and separated in 3 subsets: 50 samples for the training set (used in the APO and the fewshot algorithms), 50 samples for the validation set (used for the selection of prompts at each generation) and 200 were used as test set to allow a meaningful comparison of different prompts while limiting the risks of overfitting on other subsets during the optimzation process.
Those numbers were chosen as a tradeoff between the budget allowed to the optimization process and the typical size of the datasets which can be obtained in real life optimization tasks. Most results displayed in this paper use the ETHOS dataset.
@@
The GAAPO model had the following fixed strategy weights:
  • Random mutator: 0.4
  • APO: 0.2
  • OPRO: 0.2
  • Fewshot: 0.1
  • Crossover: 0.1
@@
Opttimization of the selection process
@@
  • Complete evaluation, which tests every prompt against every sample, requires 2,500 LLM calls (50 prompts×50 samples50 $\text{ prompts}\times 50\text{ samples}50$ prompts × 50 samples), providing exhaustive but computationally intensive evaluation.
@@
  • The bandit selection method [29] provides the most efficient tradeoff [12], evaluating only 20 prompts on 15 samples over 5 iterations. This approach requires approximately 1,500 LLM calls (20 prompts×15 samples×5 iterations20 $\text{ prompts}\times 15\text{ samples}\times 5\text{ iterations}20$ prompts × 15 samples × 5 iterations), achieving a 40% reduction in computational cost compared to complete evaluation. While this method samples less extensively, it leverages statistical efficiency to identify high-performing prompts.
@@ 5.1 Comparison with baselines
The experimental results demonstrate the effectiveness of our proposed GAAPO (Genetic Algorithm Assisted Prompt Optimization) approach on the ETHOS multilabel hate speech classification task. Figure 4 illustrates the validation performance across different prompt optimization strategies over multiple iterations, while Table 2 presents the final test and validation scores. Additionnally, obtained prompts are presented in annex Prompt obtained.
@@
A critical analysis of test and validation scores reveals an important phenomenon common to genetic algorithms: selection bias. This is particularly evident in APO’s performance, which achieves a perfect test score (1.0) at one point during training while maintaining a much lower validation score (0.40). This extreme disparity illustrates how genetic algorithms can inadvertently optimize for specific test set characteristics rather than general problem-solving capabilities. GAAPO mitigates this selection bias through its diverse strategy portfolio, resulting in more balanced performance between test (0.60) and validation (0.46) scores, suggesting better generalization.
@@ 5.4 Prompt generators comparison
[…]
@@ 5.6 Applications on other datasets
However, the results also reveal that not all tasks benefit equally from prompt optimization. The MMLU-Pro Business dataset, with its high initialization score of 0.72, shows minimal room for improvement, with our method and the Mutator achieving only marginal gains (0.73 and 0.735 respectively). This suggests that some tasks may already be well-aligned with LLMs’ base capabilities, limiting the potential impact of prompt optimization.
The GPQA dataset presents another interesting case where all optimization methods, including ours, achieve similar modest improvements (from 0.38 to 0.43), indicating that some tasks may have inherent complexity barriers that prompt optimization alone cannot overcome.
These findings suggest that the effectiveness of prompt optimization is highly task-dependent, with our method showing particular strength in tasks where there is significant room for improvement from the baseline performance.

### v2

@@ 3 Data
  • The MMLU-Pro (Massive Multitask Language Understanding Professional)[24] dataset extends the famous MMLU [25] dataset by complexifying it to a professional level, with our focus on two key subcategories. The Engineering subcategory evaluates technical understanding across various engineering disciplines, testing knowledge of fundamental principles, technical specifications, and complex problem-solving approaches encountered in professional practice. The Business subcategory assesses comprehension of management principles, corporate strategy, financial decision-making, and organizational behavior through practical business scenarios. Repartition of MMLU-Pro dataset in subcategories is detailed in annex 6.
@@ 4.1.1 GAAPO: Genetic Algorithmic Applied to Prompt Optimization
The genetic algorithm framework incorporates multiple prompt generation methods, each implementing distinct optimization strategies as detailed in sections 4.1.2 and 4.1.3. These methods, summarized in Table 1, represent diverse approaches to prompt optimization, each with its own strengths and limitations. The hybrid nature of GAAPO leverages this diversity by combining these complementary strategies within a single optimization framework. This integration enables the algorithm to capitalize on the advantages of each method while mitigating their individual limitations through iterative application of varied optimization approaches. The synergistic combination of these methods allows for more comprehensive exploration of the prompt space than would be possible with any single strategy. Examples of generated prompts on a real-world task are presented in annex LABEL:annex:prompt_obtained.
@@ 4.1.3 Generation methods: random evolutions
Each mutation creates a new variant of the original prompt (examples are presented in annex 6), potentially discovering more effective formulations. Like genetic mutations in nature, these modifications can range from subtle adjustments to significant transformations, allowing for both local and global exploration of the prompt space. This random but structured approach enables the discovery of novel prompt variations that might not be obvious through deterministic methods.
@@
4.1.4 Examples
Table 4.1.4 presents prompts obtained by several methodologies on the ETHOS multilabel dataset after 10 generations. The original prompt from where all optimization trajetories started is:
A message from a user, your goal is to determine if this message is a hate speech or not: user_message. Possible class for the hate speech are: violence, directed_vs_generalized, gender, race, national_origin, disability, religion, sexual_orientation.
Method Obtained Prompt
APO Analyze the given message to determine if it contains hate speech: user_message. Definition: Hate speech is defined as language that attacks or diminishes a person or group based on attributes such as race, religion, sexual orientation, gender, disability, national origin, or other protected characteristics. Your task is to: - Provide a detailed reasoning justifying your classification. - Categorize the message into the most relevant categories from the following list. Each category should be clearly justified in your reasoning: + sexual_orientation: Attacks or offensive language targeting individuals based on their sexual orientation or gender identity. + religion: Language that attacks or demeans individuals based on their religious beliefs or affiliation. + [... definitions of other categories...]
OPRO Analyze the message to determine if it is hate speech: user_message. Consider the following factors in your analysis: - Presence of derogatory language or slurs targeting race, gender, religion, or other protected characteristics. - Any explicit or implicit threats, calls to violence, or incitement of harm. - Discriminatory statements that demean or degrade individuals or groups. - The overall tone, intent, and context of the message. Ensure your reasoning is thorough and your solution is actionable and clear. Possible categories for hate speeches are: violence, directed_vs_generalized, gender, race, national_origin, disability, religion, sexual_orientation.
Mutator Analyze the message to determine if it is hate speech, using the following structured approach: - Identify offensive language targeting protected characteristics. - Assess intent and context. - Evaluate potential harm. - Ensure bias-free analysis adhering to hate speech policies. Provide evidence-based reasoning and actionable solutions. Examples: Racial slurs dehumanize based on race; messages promoting violence incite harm; derogatory gender terms cause marginalization. Analyze: user_message.
Ours You are a team of experts tasked with determining if a given message contains hate speech. Your team consists of three roles: **The Critic**: Your first responsibility is to carefully analyze the message for any indicators of hate speech, such as offensive language, derogatory terms, or discriminatory content. Consider the context and tone of the message. **The Problem Solver**: After the critic has provided their analysis, your role is to determine whether the message qualifies as hate speech based on the indicators identified. Consider the broader implications of labeling the message as hate speech. **The Expert Reviewer**: Your role is to review the analysis and solution provided by the previous roles. Ensure that the reasoning is thorough, the solution is accurate, and the output is consistent with the examples provided. Here is the message to analyze: user_message.
Table 2: Prompts obtained via different optimization methods on the ETHOS multilabel dataset
@@ 4.3.3 Evaluation Process
  • Initial evaluation on validation set to establish baseline performance.
  • Generational evaluation to select promising candidates. Evaluations concerning advocated results in the paper where conducted twice, due to disparities in LLM response and their probabilistic answers construction.
  • Final testing on the held-out test set to measure true generalization.
@@ 4.3.4 Metrics
ETHOS multilabel dataset
For the multi-label classification task of the ETHOS dataset, we employ strict accuracy as our evaluation metric. A prediction is considered correct if and only if the set of predicted labels exactly matches the set of true labels, regardless of their order. Formally, for a sample with true labels $Y$ and predicted labels $\hat{Y}$, the binary accuracy is defined as:
$\text{accuracy}(Y,\hat{Y})=\begin{cases}1&\text{if }Y=\hat{Y}\\ 0&\text{otherwise}\end{cases}$
where $Y$ and $\hat{Y}$ are treated as sets, meaning $\{a,b\}=\{b,a\}$. The final accuracy score is then computed as the average of these binary evaluations across all samples in the dataset.
MMLU & GPQA
For MMLU-Pro and GPQA datasets, we employ standard accuracy as our evaluation metric, where a prediction is considered correct if and only if it matches the correct answer. Formally, for a sample with true answer $y$ and predicted answer $\hat{y}$, the binary accuracy is defined as:
$\text{accuracy}(y,\hat{y})=\begin{cases}1&\text{if }y\equiv\hat{y}\\ 0&\text{otherwise}\end{cases}$
where $\equiv$ denotes semantic equivalence rather than strict string matching. This equivalence consideration was necessary as these datasets provide multiple-choice answers in a standardized format (typically including punctuation marks like commas), but the LLM sometimes generated correct answers with slight formatting variations. To address this, we implemented an LLM-based evaluation system that validates semantic correctness, ensuring that superficial differences in formatting do not impact the accuracy assessment. The reliability of this approach was verified through manual inspection of a representative sample of model outputs.
@@ 4.4 Experiments
300 samples were extracted to the orignal ETHOS dataset and separated in 3 subsets: 50 samples for the training set (used in the APO and the fewshot algorithms), 50 samples for the validation set (used for the selection of prompts at each generation) and 200 were used as test set to allow a meaningful comparison of different prompts while limiting the risks of overfitting on other subsets during the optimzation process. Those numbers were chosen as a tradeoff between the budget allowed to the optimization process and the typical size of the datasets which can be obtained in real life optimization tasks. Most results displayed in this paper use the ETHOS dataset.
@@
In the GAAPO implementation, prompt generation was distributed across strategies with the following proportions: random mutations accounted for 40% of new prompts, while APO and OPRO each contributed 20%. The remaining 20% was equally divided between few-shot learning and crossover operations, each generating 10% of new prompts.
@@
Optimization of the selection process
@@
[…]
@@
[…]
@@ 5.1 Comparison with baselines
The experimental results demonstrate the effectiveness of our proposed GAAPO (Genetic Algorithm Assisted Prompt Optimization) approach on the ETHOS multilabel hate speech classification task. Figure 4 illustrates the validation performance across different prompt optimization strategies over multiple iterations, while Table 3 presents the final test and validation scores. Additionnally, obtained prompts are presented in table 4.1.4.
@@
A critical analysis of test and validation scores reveals an important phenomenon common to genetic algorithms: selection bias. This is particularly evident in APO’s performance, where results at some iterations highlight a very high difference score between test and validation sets (culminating at 0.3 for the 2nd generation). This extreme disparity illustrates how genetic algorithms can inadvertently optimize for specific test set characteristics rather than general problem-solving capabilities. GAAPO mitigates this selection bias through its diverse strategy portfolio, resulting in more balanced performance between test (0.60) and validation (0.46) scores, suggesting better generalization.
@@
Moreover, we can see a difference in the original score of the models (at iteration 0, all scores should be identicals). However, due to the disparity in LLM performance and their probabilistic caracters, results are not always exactly consistent across time.
@@ 5.4 Prompt generators comparison
  • Strategy Effectiveness and Stability: Few-shot learning demonstrates superior performance (median $\sim$0.57) with consistent results, as shown by its compact boxplot and positive improvement scores in early generations. This aligns with existing literature [6], highlighting the value of example-based learning. OPRO maintains strong and stable performance (median $\sim$0.55), though its evolution plot shows diminishing improvements over generations. role_assignment and concise_optimization show reliable performance with tight distributions, but their improvement potential decreases in later generations.
@@ 5.6 Applications on other datasets
However, the results also reveal that not all tasks benefit equally from prompt optimization. The MMLU-Pro Business dataset, with its high initialization score of 0.72, shows minimal room for improvement, with our method and the Mutator achieving only marginal gains (0.73 and 0.735 respectively). This suggests that some tasks may already be well-aligned with LLMs’ base capabilities, limiting the potential impact of prompt optimization. The GPQA dataset presents another interesting case where all optimization methods, including ours, achieve similar modest improvements (from 0.38 to 0.43), indicating that some tasks may have inherent complexity barriers that prompt optimization alone cannot overcome.
The varying effectiveness of prompt optimization across tasks can be attributed to multiple underlying factors. First, the overlap between an LLM’s training data and the target dataset can create a ceiling effect - if similar examples were present in the training corpus, the model may already demonstrate near-optimal performance with simple prompts. Second, task-specific characteristics such as domain specificity and reasoning complexity influence the optimization potential; technical domains often benefit more from structured prompting than general knowledge tasks. Third, the nature of the required output (e.g., multiple-choice vs. multi-label classification) affects the scope for improvement through prompt engineering. Finally, the fundamental alignment between the task’s requirements and the model’s learned representations determines whether performance limitations can be addressed through prompt optimization alone or require more substantial interventions such as fine-tuning.
@@ 6 Conclusion
Acknowledgments
We extend our sincere appreciation to Joël Belafa, Florence Armstrong and Charles-Albert Lehalle whose insightful comments and suggestions substantially enhanced this work. Their expertise and careful attention to detail helped us improve various aspects of the paper, from theoretical foundations to experimental validation. Their feedback not only improved the current work but also provided valuable directions for future research in this area.